# Beyond Fact Retrieval: Episodic Memory for RAG with Generative Semantic Workspaces

### v1

@@ Appendix G Related work on Memory Augmentation for LLMs
While these structured approaches offer advantages over standard RAG by preserving more relational or hierarchical context and enabling more sophisticated information integration (like multi-hop reasoning or global summarization), they still face challenges when viewed through the lens of episodic memory (huet_episodic_2025). Graph-based methods like GraphRAG(Edge2025GraphRAG), LightRAG(Guo2024LightRAG), HippoRAG(JimenezGutierrez2025HippoRAG; gutierrez_hipporag_2025) and RAPTOR(sarthi_raptor_2024) suffer from two broad limitations. First, they lack mechanisms to track entity state/role changes across time—they represent entities as static nodes without modeling how attributes or relationships evolve throughout a narrative. Second, they provide no specific framework to ground the evolving narrative in space and time, making it difficult to represent sequential developments or causal relationships. These methods typically represent semantic relationships or summarize community structures within a potentially static corpus, but they are not explicitly designed to model the temporal flow of events within a single narrative or to meticulously track the dynamic changes in entity states and roles as the narrative unfolds sequentially. Their structure captures connections, but not necessarily the chronological progression and state transitions required for recalling specific episodes.
Other research efforts have targeted episodic memory more directly. For instance, Larimar (das_larimar_2024) proposes modifications to the LLM’s attention mechanism, while EM-LLM (fountashuman) introduces specific memory components integrated with openweight models. While promising, these approaches often require fundamental changes to the LLM architecture or are designed specifically for openweight models, limiting their applicability. In contrast, our GSW framework is proposed as a plug-and-play episodic memory module compatible with any underlying LLM (including closed-source models like GPT-4o via API) and, critically, requires no specialized training or fine-tuning of model parameters, relying instead on the LLM’s capabilities for its operator and reconciliation functions.
Appendix H Computational Costs and Resources for Building the GSW
The primary computational costs for the Generative Semantic Workspace (GSW) framework are associated with its initial, one-time indexing process. To index the 200 chapters of the Epbench dataset, the total expense is approximately \$15 when utilizing GPT-4o. This cost covers all stages of GSW construction, including the generation of operator representations, reconciliation, and the creation of entity-specific summaries. By leveraging parallel calls to the OpenAI API, managed via the Bespoke Curator library (bespoke_curator), this entire indexing task for 200 chapters can be completed in roughly 1 hour. Alternatively, the OpenAI Batch API can be used to reduce costs, with indexing taking hours.
@@ Appendix H Computational Costs and Resources for Building the GSW
Appendix I Related Computational Models of Semantics
Semantic representation frameworks have a rich history in NLP, yet as we explore below, their design choices create inherent limitations for tracking evolving actor states and relationships—a critical requirement for episodic memory. Among the most influential frameworks are PropBank (propbank) and FrameNet (framenet), which attempt to define correspondences between (a) the syntactic “realizations” of semantics explicit within language structure, and (b) finite, discrete sets of semantic “roles” (levin). These approaches rely heavily on manually-annotated lexicon ontologies developed by expert linguists. While valuable for understanding individual sentences, they were not designed for the dynamic, interconnected tracking that episodic memory demands. Below, we detail these frameworks and their limitations for serving as memory systems:
PropBank: PropBank utilized a bottom-up approach: (1) Dependency Parse Trees (dpt) were applied to a large text corpus to distill shared syntactic patterns (“Framesets”) specific to each verb (a process known as “lexical sampling”). (2) For each Frameset, the corresponding sentences were manually annotated with an enumerated set of arguments ARG:0,…, ARG:N. These arguments were later associated to verb-specific definitions using VerbNet (verbnet). The semantic roles are identified as corresponding spans within the sentence (commonly a NP, NNP subtree in the dependency parsing). For example, the sentence (A):
@@ Appendix I Related Computational Models of Semantics
Perhaps the greatest benefit of PropBank was that its syntactic “grounding” made it possible for rule-based and early ML models (rml; srlbert) to learn the task of distilling the semantics given a sentence, albeit within the confines of a limited ontology of >3000>3000 verbs and >4000>4000 Framesets
Event Databases: PropBank evolved in several directions, including efforts to unify it with related semantic lexicon such as VerbNet and FrameNet (semlink; semlink2), or augment it via the DWD overlay (dwdoverlay) to WikiData (wikidata). The latter of these efforts now manifests as “Event” databases (eventreview; xiang2019survey) such as the ACE (ace) and ERE (ere) datasets, and led to the DARPA initiative of Event identification/extraction challenges. Events are best motivated by their related identification tasks: Given a sentence, identify the event(s) – from a set of hundreds of events in a pre-annotated schema (glen; eventextract; text2event) – that the sentence is referring to. For example, (A) would be annotated with the Capture event.
@@
FrameNet (thousands of frames and tens of thousands of FE) is a substantially larger and more comprehensive ontology (framenetcompare) compared to Propbank. When originally constructed, automated systems could not effectively identify the frames implied by a sentence; today, however, Transformer models (attention; bert; fst) have demonstrated success at accurately modeling the sentence-to-frame mapping.
@@
More recent work  (conceptnet) has attempted to assemble a comparatively larger (and less stringent) open-schema semantic ontology of concepts using game-play based crowd-sourcing techniques (verbosity). However, such efforts to scale manual annotation ultimately do not address how a complete ontology can be constructed. Event Graph Models (EGM) (tegm) generate event networks to describe the dynamics of events in a text corpus, often using a combination of submodules such as Coreference Resolution (coref), Named Entity Recognition (NER) (ner) and Semantic Role Labeling (srl). Extensions (future) generate the most likely event template sequences. These methods rely on predefined event schema to enumerate the set of possible events. However, while EGMs both track the evolution of semantics across sentences and offer an unsupervised approach to extending existing ontologies, these often marginalize across individual contexts in the training corpus, and generate the most likely event schema that follows a current event schema network. As a result, these works have yet to design methods to track the semantics across a specific document. The GSW, particularly through the operator, is designed to overcome these challenges by generating actor-centric, evolving semantic maps that are not constrained by predefined, static lexicons and can capture the nuances of unfolding situations.To demonstrate the GSW Operator’s effectiveness in producing these comprehensive semantic maps from complex, real-world text, we conducted a qualitative human evaluation.
@@
To empirically validate the Operator’s capability in generating these comprehensive semantic maps, particularly its proficiency in interpreting narrative-rich texts where actor, roles and states undergo significant evolution, we leveraged news reports, as they are a popular resource for sampling semantics-rich stories that belong to universally recognized situation patterns. We query GDELT (gdelt), a Jigsaw-powered news-indexing platform, with Situation identifiers to retrieve a small set of situation-conditioned en_US articles. Table  14 presents statistics about the data. These situations were manually selected as an initial seed set – similar to FrameNet’s early versions containing few frames (fillmore) – to assess the validity of the GSW framework. Situation-specific seeds are assembled using a bootstrapped method that invokes FrameNet: (a) Frames are linked using subframe and precedence relationships to create weakly connected components; (b) Headers/labels of the frames in each component form the seed search phrases. We evaluate our framework on situations like “Crime and Justice”, “Fire Fighting”, “Healthcare”, and “Technology Development”.

### v2

@@ G Related work on Memory Augmentation for LLMs
While these structured approaches offer advantages over standard RAG by preserving more relational or hierarchical context and enabling more sophisticated information integration (like multi-hop reasoning or global summarization), they still face challenges when viewed through the lens of episodic memory (Huet et al. 2025). Graph-based methods like GraphRAG(Edge et al. 2025), LightRAG(Guo et al. 2025), HippoRAG(Gutiérrez et al. 2025b, a) and RAPTOR(Sarthi et al. 2024) suffer from two broad limitations. First, they lack mechanisms to track entity state/role changes across time—they represent entities as static nodes without modeling how attributes or relationships evolve throughout a narrative. Second, they provide no specific framework to ground the evolving narrative in space and time, making it difficult to represent sequential developments or causal relationships. These methods typically represent semantic relationships or summarize community structures within a potentially static corpus, but they are not explicitly designed to model the temporal flow of events within a single narrative or to meticulously track the dynamic changes in entity states and roles as the narrative unfolds sequentially. Their structure captures connections, but not necessarily the chronological progression and state transitions required for recalling specific episodes.
Other research efforts have targeted episodic memory more directly. For instance, Larimar (Das et al. 2024) proposes modifications to the LLM’s attention mechanism, while EM-LLM (Fountas et al. 2024) introduces specific memory components integrated with openweight models. While promising, these approaches often require fundamental changes to the LLM architecture or are designed specifically for openweight models, limiting their applicability. In contrast, our GSW framework is proposed as a plug-and-play episodic memory module compatible with any underlying LLM (including closed-source models like GPT-4o via API) and, critically, requires no specialized training or fine-tuning of model parameters, relying instead on the LLM’s capabilities for its operator and reconciliation functions.
H Computational Costs and Resources for Building the GSW
The primary computational costs for the Generative Semantic Workspace (GSW) framework are associated with its initial, one-time indexing process. To index the 200 chapters of the Epbench dataset, the total expense is approximately \$15 when utilizing GPT-4o. This cost covers all stages of GSW construction, including the generation of operator representations, reconciliation, and the creation of entity-specific summaries. By leveraging parallel calls to the OpenAI API, managed via the Bespoke Curator library (Marten et al. 2025), this entire indexing task for 200 chapters can be completed in roughly 1 hour. Alternatively, the OpenAI Batch API can be used to reduce costs, with indexing taking hours.
@@ H Computational Costs and Resources for Building the GSW
I Related Computational Models of Semantics
Semantic representation frameworks have a rich history in NLP, yet as we explore below, their design choices create inherent limitations for tracking evolving actor states and relationships—a critical requirement for episodic memory. Among the most influential frameworks are PropBank (Kingsbury and Palmer 2002) and FrameNet (Baker et al. 1998), which attempt to define correspondences between (a) the syntactic “realizations” of semantics explicit within language structure, and (b) finite, discrete sets of semantic “roles” (Levin 1993). These approaches rely heavily on manually-annotated lexicon ontologies developed by expert linguists. While valuable for understanding individual sentences, they were not designed for the dynamic, interconnected tracking that episodic memory demands. Below, we detail these frameworks and their limitations for serving as memory systems:
PropBank: PropBank utilized a bottom-up approach: (1) Dependency Parse Trees (Nivre 2010) were applied to a large text corpus to distill shared syntactic patterns (“Framesets”) specific to each verb (a process known as “lexical sampling”). (2) For each Frameset, the corresponding sentences were manually annotated with an enumerated set of arguments ARG:0,…, ARG:N. These arguments were later associated to verb-specific definitions using VerbNet (Schuler 2005). The semantic roles are identified as corresponding spans within the sentence (commonly a NP, NNP subtree in the dependency parsing). For example, the sentence (A):
@@ I Related Computational Models of Semantics
Perhaps the greatest benefit of PropBank was that its syntactic “grounding” made it possible for rule-based and early ML models (Johansson and Nugues 2008; Shi and Lin 2019) to learn the task of distilling the semantics given a sentence, albeit within the confines of a limited ontology of >3000>3000 verbs and >4000>4000 Framesets
Event Databases: PropBank evolved in several directions, including efforts to unify it with related semantic lexicon such as VerbNet and FrameNet (Palmer 2009; Shi and Mihalcea 2005), or augment it via the DWD overlay (Spaulding et al. 2023) to WikiData (Vrandečić and Krötzsch 2014). The latter of these efforts now manifests as “Event” databases (Liu et al. 2020; Xiang and Wang 2019) such as the ACE (Doddington et al. 2004) and ERE (Aguilar et al. 2014) datasets, and led to the DARPA initiative of Event identification/extraction challenges. Events are best motivated by their related identification tasks: Given a sentence, identify the event(s) – from a set of hundreds of events in a pre-annotated schema (Zhan et al. 2023; Wadden et al. 2019; Lu et al. 2021) – that the sentence is referring to. For example, (A) would be annotated with the Capture event.
@@
FrameNet (thousands of frames and tens of thousands of FE) is a substantially larger and more comprehensive ontology (Baker 2017) compared to Propbank. When originally constructed, automated systems could not effectively identify the frames implied by a sentence; today, however, Transformer models (Vaswani et al. 2017; Devlin et al. 2019; Chanin 2023) have demonstrated success at accurately modeling the sentence-to-frame mapping.
@@
More recent work  (Speer et al. 2017) has attempted to assemble a comparatively larger (and less stringent) open-schema semantic ontology of concepts using game-play based crowd-sourcing techniques (von Ahn et al. 2006). However, such efforts to scale manual annotation ultimately do not address how a complete ontology can be constructed. Event Graph Models (EGM) (Mellor 2017) generate event networks to describe the dynamics of events in a text corpus, often using a combination of submodules such as Coreference Resolution (Ng and Cardie 2002), Named Entity Recognition (NER) (Li et al. 2020) and Semantic Role Labeling (Palmer et al. 2011). Extensions (Li et al. 2021) generate the most likely event template sequences. These methods rely on predefined event schema to enumerate the set of possible events. However, while EGMs both track the evolution of semantics across sentences and offer an unsupervised approach to extending existing ontologies, these often marginalize across individual contexts in the training corpus, and generate the most likely event schema that follows a current event schema network. As a result, these works have yet to design methods to track the semantics across a specific document. The GSW, particularly through the operator, is designed to overcome these challenges by generating actor-centric, evolving semantic maps that are not constrained by predefined, static lexicons and can capture the nuances of unfolding situations.To demonstrate the GSW Operator’s effectiveness in producing these comprehensive semantic maps from complex, real-world text, we conducted a qualitative human evaluation.
@@
To empirically validate the Operator’s capability in generating these comprehensive semantic maps, particularly its proficiency in interpreting narrative-rich texts where actor, roles and states undergo significant evolution, we leveraged news reports, as they are a popular resource for sampling semantics-rich stories that belong to universally recognized situation patterns. We query GDELT (Leetaru and Schrodt 2013), a Jigsaw-powered news-indexing platform, with Situation identifiers to retrieve a small set of situation-conditioned en_US articles. Table  14 presents statistics about the data. These situations were manually selected as an initial seed set – similar to FrameNet’s early versions containing few frames (Lowe 1997) – to assess the validity of the GSW framework. Situation-specific seeds are assembled using a bootstrapped method that invokes FrameNet: (a) Frames are linked using subframe and precedence relationships to create weakly connected components; (b) Headers/labels of the frames in each component form the seed search phrases. We evaluate our framework on situations like “Crime and Justice”, “Fire Fighting”, “Healthcare”, and “Technology Development”.